OpenStudy (anonymous):

A highway makes an angle of 6 with the horizontal. This angle is maintained for a horizontal distance of 5 miles. To the nearest hundredth of a mile, how high does the highway rise in this 5-mile section? Show the steps you use to find the distance.

OpenStudy (anonymous):

 adj = 5 angle = 6 tan(angle) = opp/adj You want to find the length (height) of the opposite side: So if tan(6) = opp/5 then opp = tan(6) * 5

OpenStudy (anonymous):

so am i supposed to find tan next?

OpenStudy (anonymous):

ok girl mutiply the lenght and the height and base and divide by 2

OpenStudy (anonymous):

She's looking for the height, not the area.

OpenStudy (anonymous):

oh well ok the base x weight right

OpenStudy (anonymous):

She doesn't have the height yet. She's looking for the height. The height is the length of the opposite side. That's what she's looking for. The length of the opposite side. The length of the opposite side = the height. The height is equal to tan(6) * 5

OpenStudy (anonymous):

wait what im so confused what do i do

OpenStudy (anonymous):

Do you know what a tangent is?

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

its one of the sides of the triangle?

OpenStudy (anonymous):

tangent is the ratio of the sides of the triangle

OpenStudy (anonymous):

do you know what im supposed to do?

OpenStudy (anonymous):

open a calculator and find the value of tan(6), then the result should be multiplied by 5

OpenStudy (anonymous):

or go to google and type tan(6) get the value. The final answer is 0.5255....m